# TLUE: A Tibetan Language Understanding Evaluation Benchmark

### v3

@@ Abstract
Large language models have made tremendous progress in recent years, but low-resource languages, like Tibetan, remain significantly underrepresented in their evaluation. Despite Tibetan being spoken by over seven million people, it has largely been neglected in the development and assessment of LLMs. To address this gap, we present a Tibetan Language Understanding Evaluation Benchmark, TLUE, which is also the first large-scale benchmark for measuring the proficiency of large language models in the Tibetan language. TLUE comprises two major components: a comprehensive multi-task understanding benchmark spanning 5 domains and 67 subdomains, and a safety benchmark encompassing 7 subdomains. Finally, we evaluate a diverse set of state-of-the-art LLMs. Experimental results demonstrate that most large language models perform below the random baseline, especially highlighting the considerable challenges they face in Tibetan language processing. TLUE provides a crucial foundation for advancing future research in Tibetan language understanding and highlights the importance of promoting greater inclusivity in the development of large language models.
@@ 1 Introduction
Large language models (LLMs) have made significant strides in natural language understanding, excelling in high-resource languages like English and Chinese through extensive datasets, advanced architectures, and benchmarks like GLUE Wang et al. (2018) and SuperGLUE Wang et al. (2019). However, many mainstream LLMs, like GPT OpenAI (2024); Achiam et al. (2023), Claude Anthropic (2024), Gemini DeepMind (2024), LlaMA Dubey et al. (2024), Qwen Team (2024) and DeepSeek DeepSeek-AI et al. (2025, 2024) largely overlook low-resource languages, which limits the inclusivity and fairness of AI systems and restricts their applicability to underserved language communities Bender et al. (2021); Huang et al. (2025). Tibetan is one of the languages included, spoken by over seven million people. It has unique grammatical features and limited annotated data of Statistics of China (2021). Despite its cultural importance, Tibetan is underrepresented in current LLM research and benchmarks An (2023); Liu et al. (2025).
Current language understanding benchmarks, including GLUE Wang et al. (2018), SuperGLUE Wang et al. (2019), and BIG-bench Srivastava et al. (2022), focus on high-resource languages, leaving low-resource languages like Tibetan without tailored evaluation frameworks, including benchmarks from China Li et al. (2023); Zhang et al. (2023). Even existing methods do not adequately capture Tibetan’s linguistic intricacies Liu et al. (2022); Lv et al. (2025). In this paper, we introduce the TLUE, a Tibetan Language Understanding Evaluation benchmark, the first large-scale benchmark for Tibetan, designed to address the unique challenges of low-resource language evaluation. Furthermore, we selected several state-of-the-art large language models for training and evaluation, and the experimental results demonstrate that TLUE effectively exposes the limitations of current models in handling Tibetan language tasks. Most LLMs show significantly lower performance on TLUE compared to benchmarks in high-resource languages, particularly struggling with tasks requiring deep linguistic understanding and domain-specific knowledge. This highlights the urgent need for more inclusive pretraining data and architecture adaptations tailored to low-resource languages like Tibetan. Our findings validate TLUE as a robust tool for evaluating and guiding the development of more equitable language models.
The main contributions of our work are summarized as follows:
@@
  • We evaluated several state-of-the-art large language models, including GPT-4O Achiam et al. (2023) and GPT-O1-mini OpenAI (2024), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5 DeepMind (2024), LlaMA-3.1 Dubey et al. (2024), Qwen-2.5 Team (2024), DeepSeek-V3 DeepSeek-AI et al. (2024) and DeepSeek-R1 DeepSeek-AI et al. (2025), to systematically assess their capabilities in both general Tibetan language understanding and safety-oriented tasks. These LLMs represent a diverse set of architectures and training strategies, enabling a broad comparison of their performance under low-resource conditions.
  • Our qualitative and quantitative analyses reveal substantial limitations in current large language models when applied to Tibetan. Most models performed significantly below expected baselines, even falling below random choice levels on certain tasks, particularly in domains requiring complex reasoning or cultural sensitivity. These findings underscore the urgent need for research into more inclusive, linguistically-aware model development tailored to low-resource languages.
@@ 2.2 Low Resource Language Evaluation
Several efforts have been made to extend LLM evaluation to low-resource languages. IndicGLUE Kakwani et al. (2020), and INDICGENBENCH Singh et al. (2024) focus on Indic languages, while projects such as AmericasNLP Mager et al. (2021) and Masakhane $\forall$ et al. (2020) target indigenous and African languages. Despite these advances, Tibetan remains significantly underrepresented in LLM evaluation.
@@ 2.4 Motivation
Based on the current progress of LLM in minority languages, we propose the TLUE, filling the gap by providing a comprehensive Tibetan evaluation dataset across multiple domains and safety-related tasks. We referred to the two benchmarks, CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), and asked ancient Tibetan experts to manually translate and verify them. In summary, the combination of Ti-MMLU and Ti-SafetyBench forms TLUE. TLUE can not only enable systematic evaluation of model performance in a challenging low-resource setting, but also introduces a dedicated safety evaluation suite for Tibetan, providing insights into model robustness and potential vulnerabilities in a low-resource language context.
@@ 3.1 Curation Process
Following this, expert human annotators, including 2 Tibetan language specialists and their team of 5 annotators, refine the dataset to preserve linguistic accuracy and classical Tibetan grammar. Annotators are compensated at an hourly rate of 28 USD, ensuring high-quality review and incentivizing skilled professionals.
@@ 3.3 Why Extend Existing Benchmarks?
3.3.1 Enabling Cross-Lingual Comparability
A translation-based benchmark provides parallel data to compare Tibetan model performance across languages, isolating linguistic effects from task knowledge.
3.3.2 Addressing Data Scarcity and Leveraging Established Designs
Due to limited Tibetan corpora, TLUE adapts reliable, pre-validated benchmarks like CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), ensuring task diversity and alignment with Tibetan’s linguistic and cultural traits.
3.3.3 Ensuring Quality through Controlled Adaptation
TLUE uses a multi-stage adaptation process (translation, filtering, alignment, annotation) to balance efficiency with linguistic precision, maintaining high benchmark standards. This structured approach provides a high-quality, scalable framework for assessing LLM performance in Tibetan, especially in low-resource settings.
@@ 4 Implementation
Our experiments cover a diverse range of LLMs, including both open-source and proprietary LLMs.
For Open-source LLM, we evaluate several open-source LLMs, including LlaMA-3.1-405B Dubey et al. (2024), LlaMA-3.1-8B Dubey et al. (2024), Qwen-2.5-72B Team (2024), DeepSeek-V3 DeepSeek-AI et al. (2024), and DeepSeek-R1 DeepSeek-AI et al. (2025). These LLMs represent a variety of architectures and parameter scales, allowing for a broad comparison of open-source advancements in Tibetan language understanding.
For Proprietary LLM, we also evaluate several proprietary LLMs, including GPT-4O OpenAI (2024), GPT-3.5-Turbo Achiam et al. (2023), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5-Flash-Latest DeepMind (2024), and O1-Mini OpenAI (2024). These models serve as strong baselines for commercial LLM performance in Tibetan tasks.
@@ 4.1 Experimental Setup
$\bullet$ We study the effect of LLM size on Tibetan language understanding, comparing different parameter variants of LlaMA-3.1 Dubey et al. (2024) and Qwen-2.5 Team (2024).
@@ 4.2.1 Evaluation Methodology
We employ zero-shot evaluation to simulate real-world conditions where models must perform tasks without task-specific examples, providing a more accurate assessment of performance in Tibetan. Appendix Figure 7 illustrates the TLUE prompt format and example queries used in evaluation. To ensure consistency, we use the default temperature settings for each model during inference, avoiding bias from manual adjustments.
@@ 4.2.2 Evaluation Metrics
We evaluate LLM performance on TLUE using two metrics: response rate and accuracy. Response Rate measures the proportion of valid responses, while Accuracy tracks the proportion of correct answers. We also compute Conditional Accuracy, which reflects the LLM’s performance when giving valid responses.
@@
These methods provide a thorough assessment of LLMs’ Tibetan language performance. For more details of Ti-MMLU and Ti-SafetyBench, please refer to Appendix B Algorithm 1 and Appendix B Algorithm 2.
@@ 5 Performance on TLUE
This section shows the performance of LLM on the TLUE benchmark. We use abbreviated forms for some of the special names or evaluation matrix names. Tables 1(a) and 1(b) show results on Ti-MMLU and Ti-SafetyBench, evaluating Tibetan language understanding and safety.
Method LLM Version Avg. STEM Human Social Other China
DA Claude 3.5-Sonnet 33.95 28.04 36.32 39.31 34.01 32.09
Gemini 1.5-Flash 30.14 25.75 29.72 35.9 29.61 29.7
DeepSeek R1 15.74 13.87 14.24 18.06 13.94 18.58
DeepSeek V3 29.51 23.57 31.97 33.65 29.92 28.44
GPT 4O 16.00 12.73 16.36 18.04 16.92 15.96
GPT 3.5-Turbo 2.11 2.40 2.18 2.20 1.88 1.87
GPT O1-mini 6.14 6.15 6.33 7.17 6.06 4.98
Qwen 2.5-32B 13.94 12.63 14.98 14.92 13.71 13.44
Qwen 2.5-72B 7.27 6.07 7.98 7.52 7.74 7.02
Qwen 2.5-7B 1.8 2.94 1.87 1.63 0.9 1.68
LlaMA 3.1-405B 25.08 23.88 24.25 25.58 27.62 24.07
LlaMA 3.1-70B 23.73 23.16 23.2 26.2 24.65 21.45
LlaMA 3.1-8B 5.47 5.48 5.56 5.99 5.46 4.86
CAA Claude 3.5-Sonnet 35.63 30.88 37.47 40.58 35.26 33.96
Gemini 1.5-Flash 31.01 26.68 30.58 36.69 30.30 30.81
DeepSeek R1 27.45 21.01 25.99 32.72 26.44 31.08
DeepSeek V3 32.16 27.03 34.58 36.26 32.00 30.94
GPT 4O 17.51 14.25 17.71 19.69 18.46 17.45
GPT 3.5-Turbo 3.42 3.82 3.35 3.68 3.09 3.16
GPT O1-mini 9.67 9.69 9.80 10.14 9.68 9.02
Qwen 2.5-32B 18.56 16.66 20.3 19.72 17.47 18.67
Qwen 2.5-72B 16.50 15.73 17.88 17.00 15.84 16.04
Qwen 2.5-7B 14.59 13.92 13.66 16.34 14.57 14.46
LlaMA 3.1-405B 25.28 24.10 24.50 25.87 27.73 24.22
LlaMA 3.1-70B 23.79 23.22 23.24 26.31 24.65 21.52
LlaMA 3.1-8B 7.44 7.95 7.54 7.38 7.41 6.92
- Random - 25.00 25.00 25.00 25.00 25.00 25.00
(a) Accuracy Performance on the Ti-MMLU
Method LLM Version Avg. OFF UB PH MH IA EM PP
DA Claude 3.5-Sonnet 50.5 32.3 56.1 63.7 61.0 31.7 52.6 65.6
Gemini 1.5-Flash 43.7 32.7 44.2 59.1 49.4 37.6 43.6 46.3
DeepSeek R1 24.3 20.5 37.2 25.8 20.4 16.9 29.2 16.7
DeepSeek V3 37.4 25.5 38.1 40.5 39.1 44.3 38.3 36.9
GPT 4O 31.1 24.0 28.8 24.6 42.1 39.7 22.4 37.6
GPT 3.5-turbo 9.4 8.6 12.3 9.7 8.7 7.3 8.5 10.6
GPT O1-mini 10.9 11.1 17.0 9.7 9.5 7.0 10.9 9.5
Qwen 2.5-72B 21.2 19.8 44.1 15.0 16.4 11.8 19.9 15.9
Qwen 2.5-32B 21.9 19.9 37.6 17.6 20.3 18.1 17.8 18.6
Qwen 2.5-7B 9.0 10.0 16.8 7.3 5.7 6.6 7.7 7.2
LlaMA 3.1-405B 43.5 36.8 31.4 46.0 52.2 50.4 44.8 46.2
LlaMA 3.1-70B 37.0 32.2 37.3 30.0 40.9 44.0 34.2 39.1
LlaMA 3.1-8B 9.9 9.7 10.9 9.4 10.1 8.8 10.5 9.9
CAA Claude 3.5-Sonnet 58.5 51.1 56.5 66.1 66.9 50.1 57.9 67.6
Gemini 1.5-Flash 49.6 44.2 44.3 60.2 56.0 44.9 51.6 51.0
DeepSeek R1 46.8 42.9 45.7 51.1 50.0 45.7 55.8 33.9
DeepSeek V3 48.3 44.3 44.9 51.1 46.4 55.6 51.8 43.6
GPT 4O 32.9 28.7 30.1 25.5 42.9 40.9 24.8 38.7
GPT 3.5-turbo 11.6 12.5 16.0 11.4 10.3 8.2 10.8 11.8
GPT O1-mini 15.1 16.3 23.1 13.5 13.4 9.9 15.4 11.7
Qwen 2.5-72B 34.1 34.8 51.6 30.9 30.7 25.5 31.3 30.6
Qwen 2.5-32B 30.6 36.0 45.2 28.5 24.4 22.6 28.6 24.9
Qwen 2.5-7B 30.2 35.2 39.3 27.1 25.9 23.2 31.2 25.8
LlaMA 3.1-405B 43.9 37.6 31.7 46.5 52.4 50.8 45.1 46.4
LlaMA 3.1-70B 37.4 32.9 37.3 30.5 41.1 44.5 35.0 39.3
LlaMA 3.1-8B 12.0 12.4 12.6 11.9 12.3 10.1 13.0 11.5
- Random - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
(b) Accuracy Performance on the Ti-SafetyBench

### v5

@@ Abstract
Large language models have made tremendous progress in recent years, but low-resource languages, like Tibetan, remain significantly underrepresented in their evaluation. Despite Tibetan being spoken by over seven million people, it has largely been neglected in the development and assessment of large language models. To address this gap, we present a Tibetan Language Understanding Evaluation Benchmark, TLUE, the first large-scale benchmark for measuring the proficiency of LLMs in the Tibetan language. TLUE comprises two major components: a comprehensive multi-task understanding benchmark spanning 5 domains and 67 subdomains, and a safety benchmark encompassing 7 subdomains. Then, we evaluate a diverse set of state-of-the-art large language models. Experimental results demonstrate that most large language models perform below the random baseline, highlighting the considerable challenges they face in Tibetan language processing. TLUE provides a crucial foundation for advancing future research in Tibetan language understanding and highlights the importance of promoting greater inclusivity in the development of large language models.
@@ 1 Introduction
Large language models (LLMs) have made significant strides in natural language understanding, excelling in high-resource languages like English and Chinese through extensive datasets, advanced architectures, and benchmarks like GLUE Wang et al. (2018) and SuperGLUE Wang et al. (2019). However, many mainstream LLMs, like GPT OpenAI (2024); Achiam et al. (2023), Claude Anthropic (2024), Gemini DeepMind (2024), LlaMA Dubey et al. (2024), Qwen Qwen Team (2024) and DeepSeek DeepSeek-AI et al. (2025, 2024) largely overlook low-resource languages, such as Tibetan, which limits the inclusivity and fairness of AI systems and restricts their applicability to underserved language communities Bender et al. (2021); Huang et al. (2025); Gao et al. (2025); Liu et al. (2025b). For Tibetan, it is one of the languages included, spoken by over seven million people, which has unique grammatical features and limited annotated data National Bureau of Statistics of China (2021). Despite its cultural importance, Tibetan is underrepresented in current LLM research and benchmarks An (2023); Liu et al. (2025a, 2024).
Current language understanding benchmarks, including GLUE Wang et al. (2018), SuperGLUE Wang et al. (2019), and BIG-bench Srivastava et al. (2022), focus on high-resource languages, leaving low-resource languages like Tibetan without tailored evaluation frameworks, including benchmarks from China Li et al. (2023); Zhang et al. (2023). Even existing methods do not adequately capture Tibetan’s linguistic intricacies Liu et al. (2022); Lv et al. (2025). So, in this paper, we introduce the TLUE, a Tibetan Language Understanding Evaluation benchmark, the first large-scale benchmark for Tibetan, designed to address the unique challenges of low-resource language evaluation. Furthermore, we selected several state-of-the-art LLMs for training and evaluation, and the experimental results demonstrate that TLUE effectively exposes the limitations of current models in handling Tibetan language tasks. Most LLMs show significantly lower performance on TLUE compared to benchmarks in high-resource languages, particularly struggling with tasks requiring deep linguistic understanding and domain-specific knowledge. This highlights the urgent need for more inclusive pretraining data and architecture adaptations tailored to low-resource languages like Tibetan. Our findings validate TLUE as a robust tool for evaluating and guiding the development of more equitable language models.
All in all, the main contributions of our work are summarized as follows:
@@
  • We evaluated several state-of-the-art LLMs, including GPT-4o Achiam et al. (2023) and GPT-O1-mini OpenAI (2024), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5 DeepMind (2024), LlaMA-3.1 Dubey et al. (2024), Qwen-2.5 Qwen Team (2024), DeepSeek-V3 DeepSeek-AI et al. (2024) and DeepSeek-R1 DeepSeek-AI et al. (2025), to systematically assess their capabilities in both general Tibetan language understanding and safety-oriented tasks. These LLMs represent a diverse set of architectures and training strategies, enabling a broad comparison of their performance under low-resource conditions.
  • Our qualitative and quantitative analyses reveal substantial limitations in current LLMs when applied to Tibetan. Most models performed significantly below expected baselines, even falling below random choice levels on certain tasks, particularly in domains requiring complex reasoning or cultural sensitivity. These findings underscore the urgent need for research into more inclusive, linguistically-aware model development tailored to low-resource languages.
@@ 2.2 Low Resource Language Evaluation
Several efforts have been made to extend LLM evaluation to low-resource languages. IndicGLUE Kakwani et al. (2020), and INDICGENBENCH Singh et al. (2024) focus on Indic languages, while projects such as AmericasNLP Mager et al. (2021) and Masakhane Nekoto et al. (2020) target indigenous and African languages. Despite these advances, Tibetan remains significantly underrepresented in LLM evaluation.
@@ 2.4 Motivation
Based on the current progress of LLM in minority languages, we propose the TLUE, filling the gap by providing a comprehensive Tibetan evaluation dataset across multiple domains and safety-related tasks. We referred to the two benchmarks, CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), and asked Tibetan language experts to manually translate and verify them. In summary, the combination of Ti-MMLU and Ti-SafetyBench forms TLUE. TLUE can not only enable systematic evaluation of model performance in a challenging low-resource setting, but also introduces a dedicated safety evaluation suite for Tibetan, providing insights into model robustness and potential vulnerabilities in a low-resource language context.
@@ 3.1 Curation Process
Following this, human annotators, including 2 Tibetan language specialists and a team of 5 additional annotators, refine the dataset to preserve linguistic accuracy and classical Tibetan grammar. Annotators are compensated at an hourly rate of 28 USD, ensuring high-quality review and incentivizing skilled professionals.
@@ 3.3 Why Extend Existing Benchmarks?
  • Enabling Cross-Lingual Comparability: a translation-based benchmark provides parallel data to compare Tibetan model performance across languages, isolating linguistic effects from task knowledge.
  • Addressing Data Scarcity and Leveraging Established Designs: due to limited Tibetan corpora, TLUE adapts reliable, pre-validated benchmarks like CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), ensuring task diversity and alignment with Tibetan’s linguistic and cultural traits.
  • Ensuring Quality through Controlled Adaptation: TLUE uses a multi-stage adaptation process (translation, filtering, alignment, annotation) to balance efficiency with linguistic precision, maintaining high benchmark standards. This structured approach provides a high-quality, scalable framework for assessing LLM performance in Tibetan, especially in low-resource settings.
@@ 4 Implementation
Our experiments cover a diverse range of LLMs, including both open-source and proprietary LLMs:
  • Open-source LLM: we evaluate several open-source LLMs, including LlaMA-3.1-405B Dubey et al. (2024), LlaMA-3.1-8B Dubey et al. (2024), Qwen-2.5-72B Qwen Team (2024), DeepSeek-V3 DeepSeek-AI et al. (2024), and DeepSeek-R1 DeepSeek-AI et al. (2025). These LLMs represent a variety of architectures and parameter scales, allowing for a broad comparison of open-source advancements in Tibetan language understanding.
  • Proprietary LLM: we also evaluate several proprietary LLMs, including GPT-4o OpenAI (2024), GPT-3.5-Turbo Achiam et al. (2023), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5-Flash-Latest DeepMind (2024), and O1-Mini OpenAI (2024). These models serve as strong baselines for commercial LLM performance in Tibetan tasks.
@@ 4.1 Experimental Setup
$\bullet$ We study the effect of LLM size on Tibetan language understanding, comparing different parameter variants of LlaMA-3.1 Dubey et al. (2024) and Qwen-2.5 Qwen Team (2024).
@@ 4.2.1 Evaluation Methodology
We employ zero-shot evaluation to simulate real-world conditions where LLMs must perform tasks without task-specific examples, providing a more accurate assessment of performance in Tibetan. Appendix Figure 7 illustrates the TLUE prompt format and example queries used in evaluation. To ensure consistency, we use the default temperature settings for each model during inference, avoiding bias from manual adjustments.
@@ 4.2.2 Evaluation Metrics
We evaluate LLMs’ performance on TLUE using two metrics: Response Rate, Accuracy and Conditional Accuracy.
  • Response Rate measures the proportion of valid responses.
  • Accuracy tracks the proportion of correct answers.
  • Conditional Accuracy reflects the LLM’s performance when giving valid responses.
@@
These methods provide a thorough assessment of LLMs’ Tibetan language performance. For more details of Ti-MMLU and Ti-SafetyBench, please refer to Appendix B Algorithm 1 and Appendix B Algorithm 2. We use abbreviated forms for some of the special names or evaluation metric names, as shown in Appendix Table 4.
@@ 5 Performance on TLUE
Method LLM Version Ti-MMLU Ti-SafetyBench
Avg. STEM Human Social Other China Avg. OFF UB PH MH IA EM $P$
DA Claude 3.5-Sonnet 33.95 28.04 36.32 39.31 34.01 32.09 50.5 32.3 56.1 63.7 61.0 31.7 52.6 65.6
Gemini 1.5-Flash 30.14 25.75 29.72 35.9 29.61 29.7 43.7 32.7 44.2 59.1 49.4 37.6 43.6 46.3
DeepSeek R1 15.74 13.87 14.24 18.06 13.94 18.58 24.3 20.5 37.2 25.8 20.4 16.9 29.2 16.7
DeepSeek V3 29.51 23.57 31.97 33.65 29.92 28.44 37.4 25.5 38.1 40.5 39.1 44.3 38.3 36.9
GPT 4O 16.00 12.73 16.36 18.04 16.92 15.96 31.1 24.0 28.8 24.6 42.1 39.7 22.4 37.6
GPT 3.5-Turbo 2.11 2.40 2.18 2.20 1.88 1.87 9.4 8.6 12.3 9.7 8.7 7.3 8.5 10.6
GPT O1-mini 6.14 6.15 6.33 7.17 6.06 4.98 10.9 11.1 17.0 9.7 9.5 7.0 10.9 9.5
Qwen 2.5-32B 13.94 12.63 14.98 14.92 13.71 13.44 21.2 19.8 44.1 15.0 16.4 11.8 19.9 15.9
Qwen 2.5-72B 7.27 6.07 7.98 7.52 7.74 7.02 21.9 19.9 37.6 17.6 20.3 18.1 17.8 18.6
Qwen 2.5-7B 1.8 2.94 1.87 1.63 0.9 1.68 9.0 10.0 16.8 7.3 5.7 6.6 7.7 7.2
LlaMA 3.1-405B 25.08 23.88 24.25 25.58 27.62 24.07 43.5 36.8 31.4 46.0 52.2 50.4 44.8 46.2
LlaMA 3.1-70B 23.73 23.16 23.2 26.2 24.65 21.45 37.0 32.2 37.3 30.0 40.9 44.0 34.2 39.1
LlaMA 3.1-8B 5.47 5.48 5.56 5.99 5.46 4.86 9.9 9.7 10.9 9.4 10.1 8.8 10.5 9.9
CAA Claude 3.5-Sonnet 35.63 30.88 37.47 40.58 35.26 33.96 58.5 51.1 56.5 66.1 66.9 50.1 57.9 67.6
Gemini 1.5-Flash 31.01 26.68 30.58 36.69 30.30 30.81 49.6 44.2 44.3 60.2 56.0 44.9 51.6 51.0
DeepSeek R1 27.45 21.01 25.99 32.72 26.44 31.08 46.8 42.9 45.7 51.1 50.0 45.7 55.8 33.9
DeepSeek V3 32.16 27.03 34.58 36.26 32.00 30.94 48.3 44.3 44.9 51.1 46.4 55.6 51.8 43.6
GPT 4O 17.51 14.25 17.71 19.69 18.46 17.45 32.9 28.7 30.1 25.5 42.9 40.9 24.8 38.7
GPT 3.5-Turbo 3.42 3.82 3.35 3.68 3.09 3.16 11.6 12.5 16.0 11.4 10.3 8.2 10.8 11.8
GPT O1-mini 9.67 9.69 9.80 10.14 9.68 9.02 15.1 16.3 23.1 13.5 13.4 9.9 15.4 11.7
Qwen 2.5-32B 18.56 16.66 20.3 19.72 17.47 18.67 34.1 34.8 51.6 30.9 30.7 25.5 31.3 30.6
Qwen 2.5-72B 16.50 15.73 17.88 17.00 15.84 16.04 30.6 36.0 45.2 28.5 24.4 22.6 28.6 24.9
Qwen 2.5-7B 14.59 13.92 13.66 16.34 14.57 14.46 30.2 35.2 39.3 27.1 25.9 23.2 31.2 25.8
LlaMA 3.1-405B 25.28 24.10 24.50 25.87 27.73 24.22 43.9 37.6 31.7 46.5 52.4 50.8 45.1 46.4
LlaMA 3.1-70B 23.79 23.22 23.24 26.31 24.65 21.52 37.4 32.9 37.3 30.5 41.1 44.5 35.0 39.3
LlaMA 3.1-8B 7.44 7.95 7.54 7.38 7.41 6.92 12.0 12.4 12.6 11.9 12.3 10.1 13.0 11.5
- Random - 25.00 25.00 25.00 25.00 25.00 25.00 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4